Calculate the amount of heat needed to boil 116.g of acetic acid (HCH3CO2), beginning from a temperature of 63.5°C. Round your answer to 3 significant digits. Also, be sure your answer contains a unit symbol.
Answer #1

We have Q = m x Cp x dT

For acetic acid boiling point is 118.1 oC

Here, m = 116 g, Cp = 2.043 J/g.K

dT = 118.1 – 63.5 = 54.6 oC = 54.6 + 273 = 327.6 K

Q = 116 x 2.043 x 327.6 = 77.6 kJ